Telescope Bootstrap Startup: 10 September 2012 UT

Observer: JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: KNewton (LBTO)
Instrument Support: JMorris (LBTO)

Summary

We were able to open briefly (01:20 hour), but had a bunch of problems ranging from missing leap seconds to jet kernel crash. We were able to find stars, and to verify guiding and collimation working at RFBG.

Humidity and/or clouds consumed the remainder of the night.

Vidhya got in some closed dome testing with LBTI-AO.

Preparations

We've now got the proper patched version of Build 2012F running. That has solved several of our software problems from last night.

Kevin was able to reset the Left Side Vent Door.

SX M2 hexapod position for retro is X=+2.53 Y=+1.65 Z=+4.3 RX=95 RY=207

Details

The beginning of the night was used for some additional adaptive secondary calibration.

03:42 Sky is starting to clear off. Humidity down to 80%. Open vent doors midway.

04:00 Start Mirror Ventilation - Chamber=12 T=8.7 D=6 Wind 5m/sec from SW.

04:20 restart Upper Treehouse Ventilation, open left side Vent Door, inspected level 5 for general dryness.

04:30 Authorize as IRTC @ RFBG . Retro is still on the SX side.

Open

04:30 Open at last, but we are having trouble connecting to DX M2.

Ignore shell RIP, but then PCS has a problem.
Mon Sep 10 04:38:31.895 2012 Preset alert for right IRTC mode ACQUIRE: PCS failed StartTrackingRotator on side right because ERROR: slewtotrack ignored,tracking polynomials are being recv'd but aren't usable: P t0 is too far in the past. 1.108645e+00 seconds.

Michele fixing our missing leap second.

And some more problems restarting MCSPU.

05:02 Authorize at LBTI so Vidhya can possibly see the sky. But apparently you can't do that when MCSPU is down.

Dan is helping is restart jet after it crashed.

05:15 Chamber is down to ambient temp of 9degC, glass is 12 degC.

05:21 restart OSS, MCS, PCS, IIF

05:28 ACQUIRE preset M5_0835

05:32 restart GCS without fake stars

05:34 Alfio, Enrico, Runa have revived adsecdx which was stuck in a strange state complaining about a hard disk problem.

05:43 We found the M5 star after a few minutes of searching IE=16.5 CA=-81.8

05:43 ACQUIRE preset to pointing star - adjusted Yglobal on M2 to -2.0 mm to remove the coma

05:47 ACTIVE preset, guiding is OK, seeing is 0.9 to 1.1 arcsec

Close

05:54 Close for rapidly rising humidity

LBTI closed-dome tests with Vidhya and Alfio

05:55 Authorize as LBTI on both sides

Vidhya gets an error that her authorize doesn't match instrument. This persists even after we have authorized LBTI.

05:58 restart IIF

07:00 Reconfigure DX swing arms for LBC

07:30 soft reboot lbccontrol

07:35 turn on LBC

07:43 Vidhya calls it a night.

07:44 Authorize LBTI on both sides (again, with DX M3 out of position) - DX M3 fails to move -

07:45 Authorize LBTI on both sides (again, with DX M3 out of position) - DX M3 moves correctly

07:47 Authorize LUCI on both sides (with M3s both at LBTI) - Only DX M3 moves to the SX position +26

07:49 Authorize LUCI on both sides (with SX At LBTI and DX at LUCI) - Only DX M3 moves to the correct position -26

It seems that the M3 selector can move to the correct position, the position for the opposite side, or not move at all.

07:53 Authorize LBC on DX side

08:00 closed dome LBC track preset on the right side. - succeeds.

08:02 Humidity is hovering around 91%, but we have a cap cloud sitting over top of us.

08:06 DOFPIA, /REDONLY

08:21 Turn off mirror ventilation as humidity is up to 94%. Glass temps are 10 degC.

08:25 Turn off LBC.

09:25 98% humidity, 8 degC, heavy fog

-- JohnHill - 10 Sep 2012
Topic revision: r3 - 10 Sep 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback